OpenStudy (anonymous):

Evaluate the limit: \[\lim_{x\to1}\frac{\sqrt{x+8}-3}{x-1}\] So I know I need to somehow cancel the \(x-1\) or otherwise make the denominator non-zero, since both numerator and denominator is zero now. I'm not sure what operations I can apply though.

OpenStudy (anonymous):

id say start by rationalizing num

OpenStudy (anonymous):

do u know how to do that?

OpenStudy (anonymous):

Yes, \[(\sqrt{x+8}-3)(\sqrt{x+8}-3)\] gives me \(x-1\). However, the denominator becomes \(-1-3 x+x \sqrt{8+x}\).

OpenStudy (anonymous):

emm...be carefull :)\[\frac{\sqrt{x+8}-3}{x-1}\times \frac{\sqrt{x+8}+3}{\sqrt{x+8}+3}\]

OpenStudy (anonymous):

Sorry, that was a typo.

OpenStudy (anonymous):

so u got\[\frac{\sqrt{x+8}-3}{x-1}\times \frac{\sqrt{x+8}+3}{\sqrt{x+8}+3}=\frac{x-1}{(x-1)(\sqrt{x+8}+3)}\]

OpenStudy (anonymous):

Oh, why did I ever combine the denominator and not see the factor... Thanks!

OpenStudy (anonymous):

no problem :)